OpenStudy (anonymous):

i dont know how to write this so ill try my best. 3with the square root sign under the square root is 72x^8 and i have to simplify it

OpenStudy (anonymous):

\[\sqrt[3]{72x^8}\]

OpenStudy (anonymous):

\[\sqrt[4]{486x ^{12}y ^{14}}\]

OpenStudy (anonymous):

i solved the first one..

OpenStudy (radar):

\[\sqrt{8\times9 \times x ^{8}}\] \[\sqrt{8}\times \sqrt{9}\times \sqrt{x ^{6}}\times \sqrt{x ^{2}}\] Extracting the cube Please note that there should be a tiny 3 in the radical. \[2x ^{^{2}}\sqrt[3]{9}\times \sqrt[3]{x ^{2}}\]